OpenStudy (anonymous):

Find the inverse Laplace transform of F(s) = (2s^2 + 20s + 36)/(s^3 + 6s^2 + 9s). I partially decomposed the fractions, and I don't really know where to go from there?

OpenStudy (anonymous):

What did you get from the partiall decomposition?

OpenStudy (anonymous):

one sec, I'm typing it

OpenStudy (anonymous):

\[F(s) = (2s ^{2} + 20s + 36) [1/(9s) - 1/([9(s+3)]) - 1/[3(s+3)^2]\]

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

I'm lost. I'm assuming I then distribute the first polynomial and then Laplace each thing, but they don't seem to fit any basic transformations.

OpenStudy (anonymous):

mmm wait

OpenStudy (anonymous):

I think you can simplify the expression before the partial fractions.

OpenStudy (anonymous):

To something like this: \[\frac{2}{s} \frac{s^2+6s+9+14s+27}{s^2+6s+9}\]

OpenStudy (anonymous):

or\[\frac{2}{s}\frac{14s+27}{s^2+6s+9}\]

OpenStudy (anonymous):

hmm... that seemed to help. thanks!

OpenStudy (anonymous):

You're welcome